OpenStudy (anonymous):

f(x)=-2(x-2)^3(x-3)^2(x+1) find x-intercept

jimthompson5910 (jim_thompson5910):

plug in f(x) = 0 and solve for x f(x)=-2(x-2)^3(x-3)^2(x+1) 0=-2(x-2)^3(x-3)^2(x+1) -2(x-2)^3(x-3)^2(x+1) = 0 (x-2)^3 = 0 or (x-3)^2 = 0 or x+1 = 0 I'll let you finish

OpenStudy (anonymous):

well the (x-2)^3 is what i don't know what to do with

jimthompson5910 (jim_thompson5910):

you take the cube root of both sides

jimthompson5910 (jim_thompson5910):

the cube root of (x-2)^3 is simply x-2 the cube root of 0 is 0

jimthompson5910 (jim_thompson5910):

after taking the cube root, you go from (x-2)^3 = 0 to x-2 = 0

OpenStudy (anonymous):

do i foil the (x-3) for (x-3)^2 then?

jimthompson5910 (jim_thompson5910):

no if (x-3)^2 = 0, then x-3 = 0 I took the square root of both sides

OpenStudy (anonymous):

oh okay. thank you!

jimthompson5910 (jim_thompson5910):

np